# Self-Introspective Decoding: Alleviating Hallucinations for Large Vision-Language Models

### v2

@@ 6 Ablation Analyses
In this section, we conduct ablation analyses on the computation efficiency and hyperparameter sensitivity. More analyses about Token Selection Strategies Analyses, Larger-scale Backbone and Other Decoding Strategies are illustrated in Appendix A.6.
@@
Computation Efficiency. One primary concern of hallucination alleviation decoding methods is the computational burden. We evaluate inference time (in seconds) and peak GPU memory (in MB) on the LLaVA-1.5 7B under the POPE adversarial setting, as shown in Table 6. Contrastive Decoding (CD) methods (Leng et al., 2024; Wang et al., 2024a) involve constructing distorted raw inputs, resulting in twice the inference complexity. OPERA (Huang et al., 2024) is based on beam-search decoding and maintains a set of beams to enlarge the candidate range. Additionally, roll back mechanism in the retrospection-reallocation strategy further exacerbates computational complexity. Our SID induces vision-and-text association hallucinations by pruning large-ratio attention-important tokens in the early layers, which greatly reduces the inference time of CD up to 30.
@@ 7 Conclusion and Future Work
In this paper, we firstly re-think contrastive decoding in LVLMs and empirically find that vision-and-text-agnostic input disturbances in CD do not always amplify desired hallucinations rather than induce potential uncertainty noise. To mitigate these issues, we propose a training-free decoding strategy named Self-Introspective Decoding (SID). By developing Context and Text-aware Token Selection (CT2S) strategy, SID amplifies vision-and-text association hallucinations to guide LVLMs in contrastive decoding, thereby improving faithfulness. Extensive experiments validate the effectiveness and robustness of SID.444Codes are available at https://anonymous.4open.science/r/SID-1795.
As for Future Work: 1) As the pruning ratios and layer are set manually, we consider training the external network to automatically determine optimal hyperparameters, inspired by (Chen et al., 2023b). 2) Moreover, given that SID amplifies fine-grained hallucinations, we consider leveraging the CT2S strategy to automatically construct high-quality negative instruction for robust visual instruction tuning rather than relying on expensive GPT-4 (Liu et al., 2024a; Zhao et al., 2023).
@@ Appendix A Appendix
Overview. More related work are described in Appendix A.1, including backgrounds about Large Vision-Language Models and Decoding Strategy in LLMs. The detailed experimental settings are in Appendix A.2 including Model Details and Implementation Details. Evaluation metric details, including CHAIR Evaluations, POPE Evaluations, and GPT-4 Assisted Evaluations are introduced in Appendix A.3. We demonstrate the GPT4-V Assisted Evaluation in Appendix A.4 for comprehensive evaluations. We visualize some case study from the LLaVA-Bench-in-the-Wild dataset for qualitative analysis in Appendix A.5. More ablation analyses including Quantitative Attention Scores Analyses, Token Selection Strategies Analyses, Larger-scale LVLM Backbone, and Adopting Other Decoding Strategies are illustrated in Appendix A.6
@@ A.1 More backgrounds
Large Vision-Language Models. Motivated by the success of Large Language Models (LLMs) (Touvron et al., 2023a; Bai et al., 2023a; Chiang & Li, 2023; Taori et al., 2023; Touvron et al., 2023b; Meta, 2024), recent studies have extended LLMs to multimodal regions and provided Large Vision-Language Models (LVLMs) (Liu et al., 2023a; Zhu et al., 2023; Chen et al., 2023a; Ye et al., 2023; Li et al., 2023a; Bai et al., 2023b; Li et al., 2023c; Dai & et al, 2023; Liu et al., 2024b; Bavishi et al., 2023; Chen et al., 2024c; Li et al., 2024) powered by pre-trained LLMs. LVLMs understand and generate diverse content in a more comprehensive way by integrating user instruction and vision inputs. LLaVA (Liu et al., 2023a) connects open-set vision encoder with LLMs (i.e., Vicuna (Chiang & Li, 2023)) by instruction tuning with elaborated language-image instruction-following data. Then, LLaVA-1.5 (Liu et al., 2024b) develops the vision-language connector that is data-efficient and powerful for better multimodal understanding. Shikra (Chen et al., 2023a) further incorporates grounding data and trains the model to understand the grounding knowledge in the given images. BLIP-2, InstructBLIP, and MiniGPT-4 (Li et al., 2023c; Dai & et al, 2023; Zhu et al., 2023) introduce a learnable querying transformer to fusion multimodal features and largely reduce image tokens. Fuyu (Bavishi et al., 2023) proposes a vanilla decoder-only architecture without the vision encoder and adapter that makes it easier to understand, scale, and deploy. InternVL (Chen et al., 2024c) three simple but effective improvements, including a strong vision encoder, dynamic high-resolution, and high-quality bilingual dataset. Recently, built on SOTA open-source LLaMA 3 (Meta, 2024) and increasing the input vision resolution to 4 more pixels, LLaVA-NeXT (Li et al., 2024) exhibits excellent multimodal capabilities. Despite the impressive results, all of the above LVLMs suffer from serious hallucination problems, and we mainly conduct experiments on advanced LVLMs, including InstructBLIP, Shikra, LLaVA-1.5, and LLaVA-NeXT.
@@
Decoding Strategy in LLMs. Selecting decoding strategies in language models is crucial, as it determines how models generate text. Greedy decoding selects the highest probability next token at each step but might lead to less varied text. Beam search (Graves, 2012) is an accumulated-score-based decoding strategy. It maintains a set of beams to enlarge the candidate range and finally selects the best one in beams, which is more sophisticated than greedy decoding. Sampling decoding generates the next words by randomly selecting from the output distribution, where Top-k sampling (Fan et al., 2018) samples from Top-k likely tokens (Fan et al., 2018) and brings diversity but sometimes induces less coherent outputs. Top-p (Nucleus) sampling (Holtzman et al., 2020) improves Top-k sampling that considers the dynamic number of words that reach the probability p, achieving a balance between randomness and relevance. Recently, to alleviate the hallucination issue, DoLa (Chuang et al., 2024) decoding emphasizes the knowledge of mature layers and downplays that of pre-mature layers. OPERA (Huang et al., 2024) is established on beam-search decoding strategy and finds the interesting phenomenon of high-probability co-occurrence between the hallucination and the knowledge aggregation patterns. OPERA penalizes ‘Over-Trust Logit’ in the beam score to alleviate aggregation patterns. Self-Introspective Decoding (SID) can be seamlessly integrated into different decoding strategies to mitigate hallucinations without sacrificing text generation quality, such as diversity, coherence, and relevance.
@@ A.2 Detailed experimental settings
Models Details. To validate the effectiveness of our SID, we conduct experiments on four representative LVLMs: InstructBLIP (Dai & et al, 2023), Shikra (Chen et al., 2023a), LLaVA-1.5 (Liu et al., 2024b), and LLaVA-NeXT (Li et al., 2024). InstructBLIP employs Q-former (Li et al., 2023c) to encode 32 image tokens to feed into LVLMs, while Shikra, LLaVA-1.5, and LLaVA-NeXT directly leverage linear projection layers as vision-language connectors to align multimodal features. Shikra and LLaVA-1.5 encode 256 and 576 image tokens to LVLMs. LLaVA-NeXT increases the input vision resolution by 4 to capture more visual details, resulting in 4 more encoded vision tokens than LLaVA-1.5. All LVLMs utilize pre-trained vision encoders like CLIP (Radford et al., 2021) vision encoder, as well as pre-trained LLMs as language decoders, such as Vicuna v1.1 (Chiang & Li, 2023), LLaMA 2 (Touvron et al., 2023b), and recently released LLaMA 3 (Meta, 2024). We provide results at the 7 Billion (B) scale, and larger-scale results are in the Appendix A.6.
@@
GPT-4 Assisted Evaluations. Besides object-existence-level hallucinations evaluated by CHAIR and POPE, GPT-4 assisted benchmark (Zhao et al., 2023) utilizes the fine-grained object-level description in the Visual Genome (VG) dataset (Krishna et al., 2017) as ground-truth and relies on the advanced GPT-4 to judge the detailed (such as positional, relational, and attribute) hallucinations and calculate Sentence-level Hallucination Ratio (SHR). With the generated sentences and manually annotated factual information, GPT-4 is prompted to evaluate whether existing hallucinations sentence by sentence. The prompt template is provided in Appendix Fig. 11. Following (Zhao et al., 2023), we utilize 200 images from the VG dataset and set max new tokens to 512, with the prompt of ‘Please describe this image in detail.’ We conduct experiments on sampling decoding strategies and representative LVLMs decoding strategies: VCD (Leng et al., 2024) and OPERA (Huang et al., 2024).
@@
A.5 Case Study.
@@
In addition to using crafted metrics (CHAIR and POPE), GPT-4/GPT4-V-aided evaluations, and MME (Fu et al., 2023) and MMBench (Liu et al., 2023b) benchmarks, we qualitatively present several case studies of SID’s hallucination alleviation ability from LLaVA-Bench-in-the-Wild dataset (Liu et al., 2023a). As illustrated in Appendix Fig. 13, 14, and 15, SID effectively mitigates hallucination in these challenging scenes by dynamically amplifying vision-and-text association hallucinations. Meanwhile, it preserves the detailness of each image.
A.6 More ABLATION ANALYSES

### v3

@@ 6 Ablation Analyses
In this section, we conduct ablation analyses on the computation efficiency and hyperparameter sensitivity. More analyses about Token Selection Strategies Analyses, Larger-scale Backbone, Other Decoding Strategies, and Visual Enhancing Decoding Strategy are in Appendix A.7.
@@
Computation Efficiency. One primary concern of hallucination alleviation decoding methods is the computational burden. We evaluate the whole dataset inference time (seconds) and peak GPU memory (MB) on the LLaVA-1.5 7B under the POPE adversarial setting, as shown in Table 6. Contrastive Decoding (CD) methods (Leng et al., 2024; Wang et al., 2024a) involve constructing distorted raw inputs, resulting in twice the inference complexity. OPERA (Huang et al., 2024) is based on beam-search decoding and maintains a set of beams to enlarge the candidate range. Additionally, roll back mechanism in the retrospection-reallocation strategy further exacerbates computational complexity. Our SID induces vision-and-text association hallucinations by pruning large-ratio attention-important tokens in the early layers, which greatly reduces the inference time of CD up to 30.
@@ 7 Conclusion and Future Work
In this paper, we firstly re-think contrastive decoding in LVLMs and empirically find that vision-and-text-agnostic input disturbances in CD do not always amplify desired hallucinations rather than induce potential uncertainty noise. To mitigate these issues, we propose a training-free decoding strategy named Self-Introspective Decoding (SID). By developing Context and Text-aware Token Selection (CT2S) strategy, SID amplifies vision-and-text association hallucinations to guide LVLMs in contrastive decoding, thereby improving faithfulness. Extensive experiments validate the effectiveness and robustness of SID.
Future Work: 1) As the pruning ratios and layer are set manually, we consider training the external network to automatically determine optimal hyperparameters, inspired by (Chen et al., 2023b). In addition, to enhance the interpretability of hallucination alleviations, we consider resorting to pre-trained analysis networks to intuitively locate spurious related vision regions. 2) Moreover, given that SID amplifies fine-grained hallucinations, we consider leveraging the CT2S strategy to automatically construct high-quality negative instruction for robust visual instruction tuning rather than relying on expensive GPT-4 (Liu et al., 2024a; Zhao et al., 2024). Note that the self-generated hallucination dataset ensures style consistency, which is crucial for preference learning (Zhao et al., 2024). Codes are available at https://github.com/huofushuo/SID.
8 Acknowledgements
This work is supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU15222621, PolyU15225023). We would like to thank Tencent AI Lab for supporting Fushuo Huo as a student researcher during his internship.
@@ Appendix A Appendix
Overview. More related work are described in Appendix A.1, including backgrounds about Large Vision-Language Models and Decoding Strategy in LLMs. The detailed experimental settings are in Appendix A.2 including Model Details and Implementation Details. Evaluation metric details, including CHAIR Evaluations, POPE Evaluations, and GPT-4 Assisted Evaluations are introduced in Appendix A.3. We demonstrate the GPT4-V Assisted Evaluation in Appendix A.4 for comprehensive evaluations. Moreover, we illustrate the results on More Benchmarks including MathVista (Lu et al., 2024), MMVet (Yu et al., 2024c), LLaVA-Bench (Liu et al., 2023a), MMMU (Yue et al., 2024a) to evaluate the complex reasoning ability in Appendix A.5. We visualize some Case Study from the LLaVA-Bench-in-the-Wild dataset for qualitative analysis in Appendix A.6. More ablation experiments and analyses including More Analyses on Other Baselines, More Analyses on Adaptive Plausibility Constraint (Eq. 3), Quantitative Attention Scores Analyses, Token Selection Strategies Analyses, Larger-scale LVLM Backbone, Adopting Other Decoding Strategies, and Visual Enhancing Decoding Strategy are illustrated in Appendix A.7
@@ A.1 More backgrounds
Large Vision-Language Models. Motivated by the success of Large Language Models (LLMs) (Touvron et al., 2023a; Bai et al., 2023a; Chiang & Li, 2023; Taori et al., 2023; Touvron et al., 2023b; Meta, 2024), with the development of multimodal learning (Huo et al., 2021b; a; 2022b; 2024c; 2022a; 2024a; 2024b; 2024d; Li & Huo, 2024), recent studies have extended LLMs to multimodal regions and provided Large Vision-Language Models (LVLMs) (Liu et al., 2023a; Zhu et al., 2023; Chen et al., 2023a; Ye et al., 2023; Li et al., 2023a; Bai et al., 2023b; Li et al., 2023c; Dai & et al, 2023; Liu et al., 2024b; Bavishi et al., 2023; Chen et al., 2024c; Li et al., 2024) powered by pre-trained LLMs. LVLMs understand and generate diverse content in a more comprehensive way by integrating user instruction and vision inputs. LLaVA (Liu et al., 2023a) connects open-set vision encoder with LLMs (i.e., Vicuna (Chiang & Li, 2023)) by instruction tuning with elaborated language-image instruction-following data. Then, LLaVA-1.5 (Liu et al., 2024b) develops the vision-language connector that is data-efficient and powerful for better multimodal understanding. Shikra (Chen et al., 2023a) further incorporates grounding data and trains the model to understand the grounding knowledge in the given images. BLIP-2, InstructBLIP, and MiniGPT-4 (Li et al., 2023c; Dai & et al, 2023; Zhu et al., 2023) introduce a learnable querying transformer to fusion multimodal features and largely reduce image tokens. Fuyu (Bavishi et al., 2023) proposes a vanilla decoder-only architecture without the vision encoder and adapter that makes it easier to understand, scale, and deploy. InternVL (Chen et al., 2024c) proposes three simple but effective improvements, including a strong vision encoder, dynamic high-resolution, and high-quality bilingual dataset. Recently, built on SOTA open-source LLaMA 3 (Meta, 2024) and increasing the input vision resolution to 4 more pixels, LLaVA-NeXT (Li et al., 2024) exhibits excellent multimodal capabilities. Some methods are also expanded to the federated version Wu et al. (2024a; ). Despite the impressive results, all of the above LVLMs suffer from serious trustworthiness issues (Liu et al., 2024f; e), especially the hallucination problem, and we mainly conduct experiments on advanced LVLMs, including InstructBLIP, Shikra, LLaVA-1.5, and LLaVA-NeXT.
@@
Decoding Strategy in LLMs. Selecting decoding strategies in language models is crucial, as it determines how models generate text. Greedy decoding selects the highest probability next token at each step but might lead to less varied text. Beam search (Graves, 2012) is an accumulated-score-based decoding strategy. It maintains a set of beams to enlarge the candidate range and finally selects the best one in beams, which is more sophisticated than greedy decoding. Sampling decoding generates the next words by randomly selecting from the output distribution, where Top-k sampling (Fan et al., 2018) samples from Top-k likely tokens (Fan et al., 2018) and brings diversity but sometimes induces less coherent outputs. Top-p (Nucleus) sampling (Holtzman et al., 2020) improves Top-k sampling that considers the dynamic number of words that reach the probability p, achieving a balance between randomness and relevance. As demonstrated in Tables 1 and 3, while greedy decoding consistently surpasses sampling decoding in terms of hallucination metrics, most open-source and closed-source LVLMs default to using sampling decoding to promote diverse and coherent chat interactions. Therefore, it is practically meaningful to consider the constraints in Eq. 3 when analyzing sampling decoding. Recently, to alleviate the hallucination issue, DoLa (Chuang et al., 2024) decoding emphasizes the knowledge of mature layers and downplays that of pre-mature layers. OPERA (Huang et al., 2024) is established on beam-search decoding strategy and finds the interesting phenomenon of high-probability co-occurrence between the hallucination and the knowledge aggregation patterns. OPERA penalizes ‘Over-Trust Logit’ in the beam score to alleviate aggregation patterns. Self-Introspective Decoding (SID) can be seamlessly integrated into different decoding strategies to mitigate hallucinations without sacrificing text generation quality, such as diversity, coherence, and relevance.
@@ A.2 Detailed experimental settings
Models Details. To validate the effectiveness of our SID, we conduct experiments on four representative LVLMs: InstructBLIP (Dai & et al, 2023), Shikra (Chen et al., 2023a), LLaVA-1.5 (Liu et al., 2024b), and LLaVA-NeXT (Li et al., 2024). InstructBLIP employs Q-former (Li et al., 2023c) to condense image tokens to 32, as a result, we are unable to visualize the dynamic token pruning process of InstructBLIP like Fig. 3 and 4. Shikra, LLaVA-1.5, and LLaVA-NeXT directly leverage linear projection layers as vision-language connectors to align multimodal features. Shikra and LLaVA-1.5 encode 256 and 576 image tokens to LVLMs. LLaVA-NeXT increases the input vision resolution by 4 to capture more visual details, resulting in 4 more encoded vision tokens than LLaVA-1.5. All LVLMs utilize pre-trained vision encoders like CLIP (Radford et al., 2021) vision encoder, as well as pre-trained LLMs as language decoders, such as Vicuna v1.1 (Chiang & Li, 2023), LLaMA 2 (Touvron et al., 2023b), and recently released LLaMA 3 (Meta, 2024). We provide results at the 7 Billion (B) scale, and larger-scale results are in the Appendix A.7.
@@
GPT-4 Assisted Evaluations. Besides object-existence-level hallucinations evaluated by CHAIR and POPE, GPT-4 assisted benchmark (Zhao et al., 2024) utilizes the fine-grained object-level description in the Visual Genome (VG) dataset (Krishna et al., 2017) as ground-truth and relies on the advanced GPT-4 to judge the detailed (such as positional, relational, and attribute) hallucinations and calculate Sentence-level Hallucination Ratio (SHR). With the generated sentences and manually annotated factual information, GPT-4 is prompted to evaluate whether existing hallucinations sentence by sentence. The prompt template is provided in Appendix Fig. 11. Following (Zhao et al., 2024), we utilize 200 images from the VG dataset and set max new tokens to 512, with the prompt of ‘Please describe this image in detail.’ We conduct experiments on sampling decoding strategies and representative LVLMs decoding strategies: VCD (Leng et al., 2024) and OPERA (Huang et al., 2024).
@@
A.5 More general benchmarks evaluation.
To further validate the general ability of self-introspective decoding strategy, we conduct experiments on complex reasoning benchmarks, including MathVista (Lu et al., 2024), MMVet (Yu et al., 2024c), LLaVA-Bench (Liu et al., 2023a), MMMU (Yue et al., 2024a). MathVista is a robust mathematical reasoning evaluation benchmark that includes a series of challenging tasks requiring detailed deep visual recognition and compositional reasoning skills. MathVista is composed of 6,141 examples derived from 31 multimodal datasets involving mathematics. We utilize the ‘testmini’ subset for testing. MMVet is a benchmark designed to evaluate large multimodal models on complex multimodal tasks. It outlines six fundamental vision-language capabilities: Recognition, Knowledge, OCR, Spatial Awareness, Language Generation, and Math, and assesses their integration across 16 emerging tasks. MMVet utilizes an LLM-based evaluator for open-ended outputs, providing consistent scoring across diverse question types and answer styles. This benchmark offers insights into the abilities of various large multimodal model paradigms and models, extending beyond mere performance ranking. LLaVA-Bench is a rigorous evaluation suite aimed at testing the visual-language alignment and instruction-following capabilities of LVLMs. Here, we utilize the LLaVA-Bench (In-the-Wild) subset, which includes 24 varied images from different domains, such as memes, paintings, and sketches, with a total of 60 questions. MMMU serves as a benchmark for assessing multimodal models on tasks that span multiple disciplines, necessitating college-level subject knowledge, and deliberate reasoning. It includes 11,500 questions across 30 subjects and 183 sub-fields, focusing on advanced perception and reasoning in domain-specific knowledge areas. We compare SID with two representative decoding methods (i.e., VCD and OPERA) on LLaVA-1.5 7B. Table 8 indicates that our SID enhances reason abilities, particularly in the LLaVA-Bench benchmark. However, VCD slightly degrades LVLM’s complex reasoning ability, as reflected in MathVista, MMVet, and MMMU benchmarks. OPERA has little gains in discrimination tasks. These results are consistent with Table 5 (MME and MMbench benchmarks), Table 3, and Table 4. We also illustrate experimental results of Existence, Count, Position, and Color subsets of the MME benchmark in Table 9, which evaluate the object-level and attribute-level hallucinations. Without the assistance of auxiliary analysis networks, our method achieves +33.4 and +18.0 compared to sampling and VCD on MME subsets, validating the effectiveness in alleviating object- and attribute-level hallucinations.
| Greedy | VCD⋆ | OPERA | SID⋆ | |
| MathVista | 27.1 | 26.3 | 27.1 | 27.4 |
| MMVet | 31.1 | 30.2 | 31.1 | 31.2 |
| LLaVA-Bench | 63.4 | 63.6 | 64.3 | 68.7 |
| MMMU | 32.6 | 31.0 | 32.6 | 33.4 |
| Object-level | Attribute-level | Total | |||
| Existence | Count | Position | Color | ||
| Sampling | 171.7 | 120.8 | 112.6 | 151.0 | 556.1 |
| Dola∗ | 173.2 | 122.4 | 115.2 | 152.8 | 563.6 |
| VCD∗ | 174.4 | 124.1 | 119.4 | 153.6 | 571.5 |
| Ours∗ | 180.5 | 130.7 | 123.8 | 154.5 | 589.5 |
[…]
| Dola⋆ | 180.1 | 127.4 | 119.3 | 154.4 | 581.2 |
| VCD⋆ | 179.5 | 128.1 | 123.8 | 155.5 | 586.9 |
| Ours⋆ | 183.9 | 132.2 | 127.8 | 155.9 | 599.8 |
A.6 Case Study.
@@
In addition to using crafted metrics (CHAIR and POPE), GPT-4/GPT4-V-aided evaluations, and MME (Fu et al., 2023) and MMBench (Liu et al., 2023b) benchmarks, we qualitatively present several case studies of SID’s hallucination alleviation ability from LLaVA-Bench-in-the-Wild dataset (Liu et al., 2023a). As illustrated in Appendix Fig. 13, 14, and 15, SID effectively mitigates hallucination in these challenging scenes by dynamically amplifying vision-and-text association hallucinations. Meanwhile, it preserves the detailness of each image. As we propose a training-free decoding method that does not rely on auxiliary analysis networks, it inherently carries over the existing weaknesses of LVLMs. Intuitive case studies, as illustrated in Appendix Fig. 13, 14, and 15, reveal that SID still generates some hallucinations, particularly in finer details such as eye color and vehicle identification specifics. These failures may be attributed to the vision encoder’s relatively limited visual perception ability. For future work, it is promising to integrate SID with InternVL (Chen et al., 2024c), which scales the vision encoder up to 6B, or consider leveraging auxiliary analysis networks like Grounding DINO (Liu et al., 2024d) or OWLv2 (Minderer et al., 2023) to mitigate LVLMs’ internal weaknesses.
A.7 More Ablation Analyses
More Analyses on Other Baselines.
As SID is a training-free plug-and-play decoding strategy that do not requires robust instruction tuning with curated datasets (Lee et al., 2022; Gunjal et al., 2024; Liu et al., 2024a; Zhao et al., 2024; Jiang et al., 2024; Yu et al., 2024b; a; Ma et al., 2024; Yue et al., 2024b) and post-hoc utilizing auxiliary analysis networks (Manakul et al., 2023; Zhou et al., 2024; Yin et al., 2023; Chen et al., 2024b; Wu et al., 2024b; Feng et al., 2024), we do not compare with above two types of methods for fairness. To further validate the robustness of SID, we compare SID with the auxiliary analysis networks-based method (i.e., LURE (Zhou et al., 2024)) and the robust instruction tuning-based method (i.e., HA-DPO (Zhao et al., 2024)) and apply the plug-and-play SID to LURE and HA-DPO on POPE, CHAIR, and MME benchmarks, following their official implementations as shown in Table 10. Experimental results indicate that SID outperforms LURE and HA-DPO in most cases. Additionally, seamlessly integrating SID with LURE and LRV can significantly enhance performance.
| POPE | CHAIRs | CHAIRi | MME | |
| Greedy | 78.6 | 55.4 | 14.2 | 805.7 |
| LURE | 78.7 | 55.1 | 14.0 | 846.2 |
| +VCD | 78.3 | 55.0 | 14.3 | 813.4 |
| +SID | 82.5 | 52.1 | 13.6 | 854.9 |
[…]
| Greedy | 83.6 | 49.6 | 14.4 | 1510.8 |
| HA-DPO | 85.4 | 44.7 | 13.6 | 1522.3 |
| +VCD | 83.0 | 46.1 | 13.8 | 1500.6 |
| +SID | 86.2 | 43.8 | 12.4 | 1532.7 |
[…]
More Analyses on Adaptive Plausibility Constraint.
| CHAIRs | CHAIRi | MME | |
| Sampling | 51.3 | 16.8 | 1471.5 |
| ICD | 48.7 | 13.9 | 1479.7 |
| w/o Eq. 3 | 52.2 | 16.6 | 1448.5 |
[…]
| w/o Eq. 3 | 51.9 | 17.2 | 1457.1 |
[…]
| w/o Eq. 3 | 46.5 | 12.2 | 1484.6 |
To further analyze the effectiveness of Eq. 3, we test contrastive decoding-based methods with and without adaptive plausibility constraint (Eq. 3) on open-end generation task (CHAIR (Rohrbach et al., 2018)) and more comprehensive MME benchmark (Fu et al., 2023). As shown in Table 11, the adaptive plausibility constraint (Eq. 3) greatly affects generated texts, and most metrics underperform vanilla sampling decoding without the adaptive plausibility constraint (Eq. 3). The proposed SID avoids potential uncertainty noise by inducing fine-grained hallucinations, greatly maintaining the performances when ablating Eq. 3.